Caspian Sea Drinks is considering the purchase of a plum juicer – the PJX5. There is no planned increase in production. The PJX5 will reduce costs by squeezing more juice from each plum and doing so in a more efficient manner. Mr. Bensen gave Derek the following information. What is the IRR of the PJX5?

a. The PJX5 will cost $2.42 million fully installed and has a 10 year life. It will be depreciated to a book value of $260,148.00 and sold for that amount in year 10.

b. The Engineering Department spent $26,614.00 researching the various juicers.

c. Portions of the plant floor have been redesigned to accommodate the juicer at a cost of $15,485.00.

d. The PJX5 will reduce operating costs by $329,148.00 per year.

e. CSD’s marginal tax rate is 27.00%.

f. CSD is 56.00% equity-financed.

g. CSD’s 15.00-year, semi-annual pay, 6.95% coupon bond sells for $1,013.00.

h. CSD’s stock currently has a market value of $20.46 and Mr. Bensen believes the market estimates that dividends will grow at 2.11% forever. Next year’s dividend is projected to be $1.48.
Answer #1

To calculate IRR, we just need project cashflow. Hence, capital structure as given in points f,g,h are irrelevant
Also, point b,c are irrelevant as both are sunk cost
Further, since at the end of 10 yr, machine shall be sold at book value, Hence, no tax impact on terminal year cash flow

Entire cashflow working and IRR calculation by hit-and-trail using excel are as below. We get 5% as project IRR

(We need to sum PV of all cashflows to zero to get correct IRR)

Year CapEx Depreciation (D) Book Value Reduction in Op Cash Flow (EBIT) FCF = EBIT (1-T)+T*(-D) with T = 27% IRR DF PV Explanation
0 -2420000 5% 1.00 -2420000 Year 0 - FCF is initial cost
1 -215985.2 2204014.8 329148.0 298594.04 5% 0.95 283398.92 Depreciation per year = (2420000-260148)/10 = 215985.2
2 -215985.2 1988029.6 329148.0 298594.04 5% 0.90 268977.06
3 -215985.2 1772044.4 329148.0 298594.04 5% 0.85 255289.11
4 -215985.2 1556059.2 329148.0 298594.04 5% 0.81 242297.73
5 -215985.2 1340074.0 329148.0 298594.04 5% 0.77 229967.46
6 -215985.2 1124088.8 329148.0 298594.04 5% 0.73 218264.67
7 -215985.2 908103.6 329148.0 298594.04 5% 0.69 207157.42
8 -215985.2 692118.4 329148.0 298594.04 5% 0.66 196615.40
9 -215985.2 476133.2 329148.0 298594.04 5% 0.62 186609.86
10 -215985.2 260148.0 329148.0 558742.04 5% 0.59 331422.39 Terminal year cash flow add sale value of machine
Total PV 0.00